OpenStudy (luhvon):

You are a real estate agent and sold a house for $89,000. If your commission is 7%, how much did you make from the sale of the house?How much did you make from the sale of the house?

OpenStudy (welshfella):

89,000 * 7 --------- 100

OpenStudy (mathstudent55):

You need to find 7% of $89,000. To find 7% of $89,000, you need to translate 7% of $89,000 into math. In math, "of" usually means "times". Also, "%" means "divide by 100". \(\sf ~~~~~~\color{red}{7\%}~~~~~~~\color{green}{of} ~~~~~~\color{blue}{$89,000}=\) \(\sf =\color{red}{\dfrac{7}{100}}~~~~~~\color{green}{\times} ~~~~~\color{blue}{$89,000}\) \(\sf =\color{red}{0.07}~~~~~~\color{green}{\times} ~~~~~\color{blue}{$89,000}\) \(=\sf 0.07 \times $89,000\) Now just do the multiplication.